# Funny Valen-Tine: Solving visual abstract reasoning problems through defining the solution distribution

### v1

@@ VI-B Funny, or not Funny
The design of the new loss function and the new decoder structure is based on a hypothesis in this paper. We hypothesize that the new regression task introduced by Funny will complete ahead of the reasoning task inherent to Valen. This premature completion of the regression task may reject further optimization or adjustment of the multi-viewpoint representations, subsequently making it difficult for these representations to optimize and adapt for the reasoning task. Therefore, we have designed the Half-split Decoder for Funny, with the aim of ensuring that the regression task introduced by Funny remains in an unfinished state from the perspective of the loss function. This ensures that the MSE loss does not decrease to zero in the early stages of optimizing the reasoning task, thereby preventing the multi-viewpoint representations from becoming stagnant.
@@ VI-C Funny for Bongard-Logo problem
So, how do we find such functions? Through Human-in-the-Loop (HITL). HITL refers to situations where human experts or users intervene and provide feedback or guidance during the operation or training of machine learning or deep learning models. By analyzing the three coexisting concepts in Bongard-Logo, we subsequently discovered that some direct data augmentation methods, such as horizontal or vertical flipping, and 90-degree, 180-degree, and 270-degree rotations, can effectively serve as the function $F(x_{i})$. Therefore, when we intend to apply the Funny method to Valen for addressing the Bongard-Logo problem, it suffices to randomly apply the aforementioned data augmentation methods to the training samples $x_{i}$ during the regular training process of Valen. It is noteworthy that, unlike in the Bongard-Logo problem, , these data augmentation methods may disrupt the key attributes in RPM problem. In other words, the Funny method tailored for the Bongard-Logo problem cannot be transferred to RPM problems, and vice versa.
@@
Given the success of the Funny method in planning the representation distribution of samples, this paper argues that it is also necessary to plan the representation of the patterns, which is $\{P_{nm}|m\in[1,M]\}_{v}$, extracted by Valen for visual reasoning problems.
@@ VIII-B Experiment on Bongard-Logo
Funny+Valen+Tine 96.0 95.2 99.3 84.6 82.9
@@ VIII-C Experiment on PGM
From the table III, it can be observed that the SBR, Tine, and Funny methods have all contributed to varying degrees of improvement in Valen’s performance on RPM problems. Valen combined with the Funny method is the first to achieve a reasoning accuracy exceeding 99.0% on the PGM problem without relying on any additional supervisory signals. Furthermore, the integration of SBR with Valen is the first approach to utilize metadata as an additional supervisory signal, achieving a reasoning accuracy of 99.4%. Additionally, SBR provides Valen with interpretability by endowing it with a progressive pattern extraction process. The SBR method presented in this paper is not the first to consider utilizing the accompanying metadata within RPM problem instances to assist the solver in reasoning. Previous works[21] have explored various approaches to leveraging metadata, but they ultimately found that forcibly imposing additional metadata learning tasks on their solvers was counterproductive. RS-Tran[28] and Triple-CFN[28] adopted an indirect manner of utilizing metadata, specifically, RS-Tran pre-trained its representation encoder using metadata, and Triple-CFN required a similar warm-start process. However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved promising results.
@@ IX Conclusion
In conclusion, this paper introduces a novel baseline model, Valen, for solving visual abstract reasoning tasks, particularly Bongard-Logo problems and RPM problems. Valen, a probabilistic-highlighting model, demonstrates remarkable performance in handling both image clustering reasoning and image progression pattern reasoning. The paper further delves into the underlying mechanisms of probability-highlighting solvers, realizing that these solvers approximate solutions to reasoning problem instances as distributions. To bridge the discrepancies between the estimated solution distribution and the true correct solution distribution, the paper introduces three methods: Tine, Funny, and SBR. The Tine method, based on adversarial learning, aims to assist Valen in estimating a solution distribution closer to the correct one. The Funny method models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. The SBR method further extends this idea by modeling the distribution of progressive patterns representation as a mixture of Gaussian distributions. Overall, the Funny, Tine, and SBR methods significantly improve Valen’s performance, providing new insights and methods for the study of visual abstract reasoning problems.

### v2

@@ VI-B Funny not funny
This Half-split Decoder can ensure that the loss of the regression task introduced by Funny does not decrease to zero in the early stage of the reasoning task of Valen, thereby preventing the multi-view representation from becoming stagnant.
@@ VI-C Funny for Bongard-Logo problem
So, how do we find such functions? Through Human-in-the-Loop (HITL). HITL refers to situations where human experts or users intervene and provide feedback or guidance during the operation or training of machine learning or deep learning models. By analyzing the three coexisting concepts in Bongard-Logo, we subsequently discovered that some direct data augmentation methods, such as horizontal or vertical flipping, and 90-degree, 180-degree, and 270-degree rotations, can effectively serve as the function $F(x_{i})$. Therefore, when we intend to apply the Funny method to Valen for addressing the Bongard-Logo problem, it suffices to randomly apply the aforementioned data augmentation methods to the training samples $x_{i}$ during the regular training process of Valen. It is noteworthy that, unlike in the Bongard-Logo problem, these data augmentation methods may disrupt the key attributes in RPM problem. In other words, the Funny method tailored for the Bongard-Logo problem cannot be transferred to RPM problems, and vice versa.
@@
This paper suggests that we should also plan the distribution of the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ extracted by Valen for visual abstract reasoning problems, so that it is closer to the distribution of the correct solution’s pattern.
@@ VIII-B Experiment on Bongard-Logo
Funny+Valen+Tine 96.0 95.2 99.6 84.6 82.9
@@ VIII-C Experiment on PGM
From the table III, it can be observed that the SBR, Tine, and Funny methods have all contributed to varying degrees of improvement in Valen’s performance on RPM problems. Valen combined with the Funny method is the first to achieve a reasoning accuracy exceeding 99.0% on the PGM problem without relying on any additional supervisory signals. Furthermore, the integration of SBR with Valen is the first approach to utilize metadata as an additional supervisory signal, achieving a reasoning accuracy of 99.4%. Additionally, SBR provides Valen with interpretability by endowing it with a progressive pattern extraction process. The SBR method presented in this paper is not the first to consider utilizing the accompanying metadata within RPM problem instances to assist the solver in reasoning. Previous works[21] have explored various approaches to leveraging metadata, but they ultimately found that forcibly imposing additional metadata learning tasks on their solvers was counterproductive. RS-Tran[28] and Triple-CFN[28] adopted an indirect manner of utilizing metadata. Specifically, RS-Tran utilizes Tranclip[28] to pre-train its representation encoder using metadata, and Triple-CFN required a similar warm-start process. However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved promising results.
@@ IX Conclusion
In conclusion, this paper introduces a novel baseline model, Valen, for solving visual abstract reasoning tasks, particularly Bongard-Logo problems and RPM problems. Valen, a probabilistic-highlighting model, demonstrates remarkable performance in handling both image clustering reasoning and image progression pattern reasoning. The paper further delves into the underlying mechanisms of probability-highlighting solvers, realizing that these solvers approximate solutions to reasoning problem instances as distributions. To bridge the discrepancies between the estimated solution distribution and the true correct solution distribution, the paper introduces three methods: Tine, Funny, and SBR. The Tine method, based on adversarial learning, aims to assist Valen in estimating a solution distribution closer to the correct one. The Funny method models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. The SBR method further extends this idea by modeling the distribution of progressive patterns representation as a mixture of Gaussian distributions. This paper demonstrates the importance of explicitly planning the distribution of solutions in solving visual abstract reasoning problems through the successful application of the SBR, Funny, Valen and Tine.